OpenStudy (anonymous):

TRIGONOMETRY a line through (2,2) and (5,-5) is parallel to another line through (o,-4) and (2,y). Find y TRIGONOMETRY a line through (2,2) and (5,-5) is parallel to another line through (o,-4) and (2,y). Find y @Mathematics

OpenStudy (anonymous):

both slope will be equal

OpenStudy (anonymous):

means (2+5)/(2-5)=(2-0)/(-4-y) solve this equation you will get right answer

OpenStudy (anonymous):

(2+5)/(2-5)=(2-0)/(-4-y) if the slope is same? so paralel means the slope always same?

OpenStudy (anonymous):

is it -22/7?